### v1

# Complementarity in Human-AI Collaboration: Concept, Sources, and Evidence

Abstract
Artificial intelligence (AI) can improve human decision-making in various application areas. Ideally, collaboration between humans and AI should lead to complementary team performance (CTP)—a level of performance that neither of them can attain individually. So far, however, CTP has rarely been observed, suggesting an insufficient understanding of the complementary constituents in human-AI collaboration that can contribute to CTP in decision-making. This work establishes a holistic theoretical foundation for understanding and developing human-AI complementarity. We conceptualize complementarity by introducing and formalizing the notion of complementarity potential and its realization. Moreover, we identify and outline sources that explain CTP. We illustrate our conceptualization by applying it in two empirical studies exploring two different sources of complementarity potential. In the first study, we focus on information asymmetry as a source and, in a real estate appraisal use case, demonstrate that humans can leverage unique contextual information to achieve CTP. In the second study, we focus on capability asymmetry as an alternative source, demonstrating how heterogeneous capabilities can help achieve CTP. Our work provides researchers with a theoretical foundation of complementarity in human-AI decision-making and demonstrates that leveraging sources of complementarity potential constitutes a viable pathway toward effective human-AI collaboration.
Keywords Human-AI Complementarity  $\cdot$ Human-AI Decision-Making  $\cdot$ Complementary Team Performance  $\cdot$ Human-AI Teams  $\cdot$ Human-AI Collaboration  $\cdot$ Complementarity Potential
1 Introduction
The increasing capabilities of artificial intelligence (AI) have paved the way for supporting human decision-making in a wide range of application domains. Examples include decision support for humans in application areas such as customer services (Vassilakopoulou et al.,, 2023), medicine (Jussupow et al.,, 2021), law (Mallari et al.,, 2020), finance (Day et al.,, 2018), and industry (Stauder and Kühl,, 2021). With AI decisions becoming increasingly accurate, there is an obvious temptation to rely on them and to automate decision tasks. However, this approach often falls short of realizing even better performance by combining the unique strengths of the individual members in a human-AI team (Seeber et al.,, 2020). The recent emergence of large language models illustrates this (Malone et al.,, 2023): While applications like ChatGPT often provide helpful, but not always correct results, a human decision-maker can collaborate with the system to, for example, override erroneous answers, in order to achieve superior task performance (Malone et al.,, 2023). Similarly, in the medical domain, both AI models and physicians are able to produce diagnoses individually. It has, however, been demonstrated that humans and AI models could make different errors (Geirhos et al.,, 2021; Steyvers et al.,, 2022). In this context, the AI model might detect patterns in large amounts of data that humans might not discover easily, while humans might excel at the causal interpretation and intuition required to interpret these patterns (Lake et al.,, 2017; Li et al.,, 2019).
This potential for complementarity has inspired researchers to investigate how humans and AI’s individual capabilities could be leveraged to achieve superior team performance instead of either performing the decision task independently. Such an outcome is defined as complementary team performance ($CTP$) (Bansal et al.,, 2021).
Various studies have demonstrated that human-AI teams could outperform human individuals, but often do not exceed the AI model’s individual performance (Bansal et al.,, 2021; Hemmer et al.,, 2021). The observation that $CTP$ is often not attained in these studies raises questions about the factors that contribute to such an outcome. Further, it illustrates that the current knowledge of how to leverage the capabilities of humans and AI for joint decision-making synergies has not yet been sufficiently developed, and that there is a need for additional concepts to foster an in-depth understanding of complementarity in human-AI decision-making. In this work, we therefore pursue the following two research questions:
RQ1: How can we model human-AI decision-making to enable a more nuanced understanding of the synergetic potential in a human-AI team?
RQ2: What factors contribute to complementary team performance?
We address these research questions by developing a conceptualization of human-AI complementarity that introduces the notion of complementarity potential ($CP$). The conceptualization comprises a formalization of this potential and a delineation of relevant sources. In detail, we argue that complementarity potential has an inherent and a collaborative component. Whereas the first encompasses effects based on different information or capabilities within the human-AI team, the second captures decision-making synergies that only emerge through human-AI interaction. We demonstrate the proposed conceptualization’s application and value in two experimental studies, in which we investigate the effect of two sources of complementarity potential on team performance—the asymmetric distributions of information and capabilities within the human-AI team. In both studies, humans collaborate with an AI model to conduct decision-making tasks. The AI model provides independent decision suggestions that humans can incorporate into their judgment to derive a final team decision. In the first study, we choose the domain of real estate valuation to investigate information asymmetry. We train an AI model to predict real estate prices based on tabular data. Humans receive its suggestions and also have access to a photograph of the real estate. They can use both to make a final team decision. In the second study, we select the context of image classification to analyze the impact of asymmetric capabilities between humans and AI. We train two AI models whose capability gaps differ from those of the human decision-maker. In both studies, we apply our conceptualization to demonstrate that both of these sources increase the inherent complementarity potential and allow $CTP$ to be attained.
To summarize, we make the following contributions to the state of knowledge in information system research: First, we conceptualize human-AI complementarity as a means to comprehensively analyze and design human-AI collaboration. Second, we scrutinize information and capability asymmetries as sources of complementarity potential. Third, we demonstrate the value and application of our concepts and the sources’ potential impact in the two behavioral experiments. This should provide IS researchers with a better understanding and with methodological support when purposefully designing human-AI collaboration in decision-making for more effective outcomes—thereby supporting the development of hybrid intelligence (Dellermann et al.,, 2019), and advocating the “AI with human” (opposed to “AI vs. human”) perspective in societal debates on the future of work (Huysman,, 2020).
In the remainder of this work, we first outline the relevant background and related work in Section 2. In Section 3, we derive the conceptualization of human-AI complementarity. In Section 4, we illustrate our conceptualization’s utility in two experimental studies. We discuss our results in Section 5, before concluding the work in Section 6. We provide the Appendix of this paper at https://github.com/ptrckhmmr/human-ai-complementarity.
2 Theoretical Foundations and Related Work
In this section, we elaborate on the key concepts and existing work on the collaboration between humans and AI, human-AI complementarity, as well as information and capability asymmetry.
2.1 Collaboration Between Humans and AI
Many terms are used to describe the collaboration between humans and AI. Common ones are human-AI team (Seeber et al.,, 2020), human-AI collaboration (Vössing et al.,, 2022), and human-AI decision-making (Lai et al.,, 2023). These are interrelated concepts that emphasize combining the capabilities of humans and AI. The notion of a human-AI team refers to an organizational setup in which AI agents [^1] Throughout this paper, we refer to agents or systems using machine learning as “Artificial Intelligence” (AI). While we acknowledge the technical distinction between AI and Machine Learning (ML) as discussed, e.g., by Kühl et al., (2022), and although we have considered the more precise reference to an “ML” agent, we adopt the use of the broader “AI” term as the prevalent terminology established in the Computer Science and Human-Computer Interaction communities. This choice reflects the contemporary linguistic trend rather than a lack of distinction between the two fields. or systems are increasingly considered team members rather than support tools for humans—since AI can perform a continuously growing number of tasks independently (Endsley,, 2023; Seeber et al.,, 2020). Human-AI collaboration is the process in which these teams work together in a synergistic manner to achieve shared goals, e.g., with the AI providing recommendations or insights, and humans guiding and refining the AI-generated outputs (Terveen,, 1995; Vössing et al.,, 2022). Human-AI decision-making specifically refers to the collaboration of humans and AI in decision-making tasks. For example, the AI could offer data-driven decision recommendations, while humans leverage their domain expertise, emotional intelligence, and ethical considerations to combine AI recommendations with their judgment to reach a final team decision (Lai et al.,, 2023).
In this article, we focus on human-AI decision-making as a key application area of human-AI collaboration. The increasing capabilities of AI have contributed to its use in a growing number of applications domains, such as medicine (McKinney et al.,, 2020), finance (Kleinberg et al.,, 2018), and customer services (Vassilakopoulou et al.,, 2023). Consequently, AI-based technologies are employed in processes and systems with varying degrees of human involvement, ranging from autonomous decision-making (Rinta-Kahila et al.,, 2022) to auxiliary support for the humans who make the final decision (Bansal et al.,, 2021; Buçinca et al.,, 2020; Lai et al.,, 2020; Liu et al.,, 2021). In this context, unintended or unfair outcomes, e.g., AI-based systems’ decisions that benefit certain individuals more than others (Kordzadeh and Ghasemaghaei,, 2022), have ignited a debate on the degree of autonomy granted to AI to ensure responsible outcomes (Mikalef et al.,, 2022). To alleviate possible detrimental effects, configurations have been suggested that keep humans in the decision-making loop (Grønsund and Aanestad,, 2020; Mikalef et al.,, 2022). Consequently, an increasing number of studies have conducted behavioral experiments to understand how humans make decisions in human-AI teams (Alufaisan et al.,, 2021; Bansal et al.,, 2021; Buçinca et al.,, 2020; Carton et al.,, 2020; Fügener et al., 2021b, ; Fügener et al.,, 2022; Lai et al.,, 2020; Liu et al.,, 2021; Malone et al.,, 2023; Reverberi et al.,, 2022; van der Waa et al.,, 2021; Zhang et al.,, 2022, 2020). In this context, a growing body of work examines how human reliance on AI decisions can be appropriately calibrated to ensure effective decision-making (Buçinca et al.,, 2020; He et al.,, 2023; Kunkel et al.,, 2019; Yu et al.,, 2019; Zhang et al.,, 2020). In application domains comprising high-stakes decisions, e.g., medicine, it is crucial for human experts to identify the AI model’s incorrect suggestions (Jussupow et al.,, 2021). Humans might be helped to judge the AI model’s decision quality by receiving information about the decision’s uncertainty (Fügener et al., 2021a, ; Zhang et al.,, 2020) or by receiving explanations that shed light on the AI model’s decision-making rationale (Bauer et al., 2023a, ).
In order to do so, explainable artificial intelligence (XAI) research has developed various approaches that aim to enable humans to understand the underlying mechanisms that contribute to an AI model’s decision (Adadi and Berrada,, 2018; Bauer et al.,, 2021). Several studies investigate how human-AI decision-making benefits from different XAI techniques. Examples of these range from feature-based techniques (Ribeiro et al.,, 2016) to example-based (van der Waa et al.,, 2021) and rule-based ones (Ribeiro et al.,, 2018). Although experiments have demonstrated the benefits of explanations (Buçinca et al.,, 2020), there is also evidence that explanations might convince humans to follow incorrect AI decisions (Bansal et al.,, 2021) or could foster humans’ propensity to delegate authority to AI by blindly approving its suggestions (Bauer et al., 2023b, ).
A closer look at quantitative studies on human-AI decision-making reveals that, in general, human performance increases when supported by high-performing AI models. In the vast majority of cases, however, the team performance remains inferior to that of the AI model performing the task alone (Hemmer et al.,, 2021; Malone et al.,, 2023). This observation is also in line with the combination of human and technical decisions that the forecasting literature has historically discussed as judgmental adjustments. In this context, several studies find that humans struggle to adjust statistical forecasts effectively (Khosrowabadi et al.,, 2022): Goodwin and Fildes, (1999) show that decision-makers often adjust highly reliable forecasts, while leaving those needing manual intervention untouched. Moreover, Lim and O’Connor, (1995) find that forecasters tend to underestimate statistical forecasts in favor of their judgments. In summary, there are many cases where human adjustments to forecasts and AI suggestions are not beneficial. However, a noteworthy scenario in which the combination of human expertise with technical forecasts could improve the forecast accuracy is in the presence of human contextual knowledge that can serve as a beneficial input for a combined forecast (Blattberg and Hoch,, 1990; Lawrence et al.,, 2006; Sanders and Ritzman,, 1995).
2.2 Human-AI Complementarity
Complementarity between humans and AI is discussed as part of several closely related paradigms: intelligence augmentation, human-machine symbiosis, and hybrid intelligence. Intelligence augmentation is defined as “enhancing and elevating human’s ability, intelligence, and performance with the help of information technology” (Zhou et al.,, 2021, p. 245). It is a form of human-AI collaboration pursuing the idea that machines use their capabilities to assist humans, not necessarily to achieve $CTP$, but to improve human objectives. Human-machine symbiosis is a paradigm that envisions deepening the collaborative connection between humans and AI. It is based on the notion of a symbiotic relationship between both and considers them as a common system rather than two separate entities with the aim to become more effective together compared to working separately (Licklider,, 1960). It also makes the assumption that both entities can offer different capabilities that can be leveraged to overcome human restrictions and to reduce the time needed to solve problems (Gerber et al.,, 2020; Jain et al.,, 2021). Hybrid intelligence pursues the idea of combining human and artificial team members in the form of a socio-technical ensemble. We refer to the work of Dellermann et al., (2019, p. 640), who define hybrid intelligence as “the ability to achieve complex goals by combining human and artificial intelligence, thereby reaching superior results to those each of them could have accomplished separately, and continuously improve by learning from each other.” In addition to realizing performance synergies on an individual level, IS research has also considered AI as a performance driver on an organizational level, when firms develop an AI capability within the organization (Mikalef and Gupta,, 2021). Nevertheless, existing studies under these labels do not provide theoretical views of human-AI complementarity. Articles by Donahue et al., (2022), Steyvers et al., (2022), and Rastogi et al., (2023) are the only works to theorize about human-AI complementarity. Donahue et al., (2022) discuss scenarios in which $CTP$ could occur by considering fairness aspects. Steyvers et al., (2022) derive a framework for combining individual decisions and different types of confidence scores from humans and AI models. Lastly, Rastogi et al., (2023) propose a taxonomy of human and AI strengths together with the notion of across- and within-instance complementarity. All three differ from our approach as they neither investigate different components of complementarity potential or effects, nor do they empirically analyze human-AI decision-making in behavioral experiments.
2.3 Information and Capability Asymmetry
In many application domains, possible sources of complementarity potential could reside in the information and capability asymmetry between humans and AI (Dellermann et al.,, 2019; Ibrahim et al.,, 2021). In this context, it is noteworthy that the mere existence of such asymmetries does not necessarily lead to performance synergies, as it could also have detrimental effects (Dougherty,, 1992).
Information Asymmetry.
While AI requires digitally available and sufficient training data to detect patterns that could subsequently be used for decision recommendations (LeCun et al.,, 2015), relevant information might not have been digitized for technical or economic reasons (Ibrahim et al.,, 2021). Contextual information might also not be available to the degree required for it to be included in the AI model training data. Humans, however, could use their expertise by also considering non-digitized information and information about rare events to create a holistic picture in decision-making situations. This might lead to performance synergies when collaborating with AI. This hypothesis is also supported by forecasting theory (Sanders and Ritzman,, 1991, 1995, 2001). Sanders and Ritzman, (1995) analyze the effects of combining statistical forecasts’ predictions with that of a human with access to contextual information, finding that such combinations could impact forecast accuracy positively. Conversely, AI models could also utilize unique contextual information when analyzing information not available to humans in particular decision-making situations. For example, in driving assistance, AI models might infer trajectory suggestions from sensor data to which the driver does not have access.
Capability Asymmetry.
Research on human teams extensively investigated team composition’s impact on performance (Horwitz,, 2005): On the one hand, asymmetric capabilities of different team members could foster performance synergies, e.g., when these lead to more comprehensive considerations that affect decision-making processes positively (Simons et al.,, 1999). On the other hand, capability asymmetry could also affect team performance negatively due to, e.g., difficulties with developing a shared understanding (Dougherty,, 1992). Similarly, there are also asymmetric capabilities in human-AI teams. Nevertheless, to date the effect of different human and AI capabilities on team performance has remained largely unexplored. Asymmetric capabilities could, e.g., emerge from different internal processing modes. Whereas AI models mostly encode statistical relationships, humans draw upon compositional mental models that encode sophisticated beliefs about the physical and social world (Lake et al.,, 2017; Rastogi et al.,, 2023). While AI models require vast amounts of data for training, humans are able to make inferences from very few data instances (Gopnik and Wellman,, 2012; Lake et al.,, 2017; Tenenbaum et al.,, 2011). In contrast, AI’s internal processing exhibits greater capability to perceive small variations in data than humans can (Findling and Wyart,, 2021). In addition to different processing modes, the nature of the “experiences” that humans and AI are exposed to could also result in asymmetric capabilities. Even though AI models are trained on large datasets, they are restricted to a limited amount of data as input. Humans, on the contrary, accumulate experiences over their entire lifetime, including information across many domains (Dellermann et al.,, 2019; Rastogi et al.,, 2023).
3 Conceptualization of Human-AI Complementarity
In this section, we first introduce the fundamental notion of human-AI complementarity, and formalize our decision-making situation as a basis for further analysis. Subsequently, we introduce and formalize the complementarity potential concept and discuss its possible sources.
3.1 The Principle of Human-AI Complementarity
We first motivate the underlying idea of complementarity which drives effective human-AI collaboration. In this work, we focus on the performance on decision-making tasks that humans and AI can conduct independently. We therefore recognize that AI contributions could, in many domains, go beyond mere (partial) decision support for humans, and could be regarded as an alternative way to solve a task independently. Real-world examples include diagnosing diseases in medicine (Goldenberg et al.,, 2019), conducting loan decisions in finance (Turiel and Aste,, 2020), and writing entire programs with AI code generation systems (Ross et al.,, 2023). However, since neither humans nor AI are perfect, their different available information or capabilities could be combined to generate superior outcomes in a human-AI team. Figure 1 illustrates the situation in a simple example for a set of binary decisions: The decision-making task comprises 25 instances with binary “right” or “wrong” outcomes. The number of incorrect decisions measures the performance. The AI makes 13 incorrect decisions, while the human errs at 15 task instances when conducting the task independently. Of all the instances, neither the AI nor the human can solve 5 on their own. Consequently, if the “human-AI team” were, for each instance, to pick the correct decision of either the AI or the human, the team could correctly solve 20 instances and would only miss the 5 that none of them can solve. In other words, human-AI collaboration opens an inherent “complementarity potential” of 8 instances compared to the individually better performing team member (here, the AI with 13 misses).
Figure 1: Illustration of the principle of human-AI complementarity.
In reality, and depending on the application context, performance could be captured by more intricate measures than just the absolute number of errors, e.g., as the precision or the recall in classification tasks (e.g., when analyzing radiology images in health care), as the mean absolute error in prediction tasks (e.g., when making sales forecasts for inventory management), or as more complex compound metrics (e.g., when weighting multiple dimensions of interest). For complementarity potential to exist, it is essential that humans and AI have different strengths and weaknesses, i.e., that they make different errors across a variety of tasks (Geirhos et al.,, 2021; Steyvers et al.,, 2022). If their capabilities could be combined adequately, the team performance in such situations would be superior to their individual performances (Rastogi et al.,, 2023). In the following, we define this potential and its components formally.
3.2 Human-AI Decision-Making Setting
Let us first define the human-AI decision-making setting illustrated in Figure 1, which is the foundation of this work: A decision task $T=\left\{(x^{(i)},\ y^{(i)})\right\}_{i}^{N}$ is a set of $N$ instances $x^{(i)}\in X$ with corresponding ground truth labels $y^{(i)}\in Y$ denoting the correct results. Each instance $x^{(i)}$ represents an individual decision. In this section, we use the term decision for both classification and prediction tasks. The ground truth, i.e., the correct decision, might not be known at the time of the decision, but can be determined and revealed later. Both a human decision-maker $H$ and a machine learning model, which we denote as $AI$, are capable of independently producing a decision for each task instance. For any given instance $x^{(i)}$, the human and the AI will independently derive decisions $\hat{y}_{H}^{(i)}$ and $\hat{y}_{AI}^{(i)}$. In a scenario where the human and the AI might collaborate in respect of their decision-making, a collaboration mechanism $I(\hat{y}_{H}^{(i)},\hat{y}_{AI}^{(i)})$ combines their decisions into a final team decision $\hat{y}_{I}^{(i)}$. We note that this decision might be different from each individual decision. Each decision’s quality is measured by its deviation (“loss”) from the ground truth—by a loss function $l$ bounded in $R^{+}$. This function serves as a generic measure of performance that could take different forms with different decision problems, e.g., an error rate used in classification tasks (like the number of wrong or unsolved task instances in Figure 1). For any instance $x^{(i)}$, losses are given by $l_{H}^{(i)}$ for the human decision, $l_{AI}^{(i)}$ for the AI decision, and $l_{I}^{(i)}$ for the combined team decision. This results in overall losses $L_{D}$ for the entire task by averaging all the available instances:
$L_{D}=\frac{1}{N}\sum_{i=1}^{N}{l_{D}^{(i)}\left({\hat{y}}_{D}^{(i)},y^{(i)}\right)}\ with\ D\ \in\ \left\{H,AI,I\right\}$ (1)
3.3 Complementarity Potential
From a decision-theoretic perspective, the vision of human-AI collaboration is to attain a superior team performance compared to the human and the AI conducting the task individually—providing the fundamental reason for forming human-AI teams (Rastogi et al.,, 2023). In our context, the human-AI team reaches complementary team performance ($CTP$) when the loss of the team is strictly smaller than that of the human and the AI individually (Bansal et al.,, 2021):
$CTP=\begin{cases}1,&L_{I}<\min(L_{H},L_{AI}),\\ 0,&\text{otherwise}.\end{cases}$ (2)
In addition to a binary task outcome, we propose the notion of complementarity potential ($CP$) to measure the discrepancy between the overall loss of the individually better performing team member $T^{\ast}\in\{H,AI\}$ with $L_{T^{\ast}}=\min\left(L_{H},L_{AI}\right)$ and perfect decisions for all instances of a task, i.e., selecting the ground truth, with an overall loss of 0:
$CP=\ L_{T^{\ast}}=min\left(L_{H},L_{AI}\right)$ (3)
In our introductory example in Figure 1, the overall loss is quantified by the number of wrong decisions. Consequently, the complementarity potential amounts to 13, which is given as the minimum number of individual errors (13 for the AI, 15 for the human). [^2] For simplification, we report $L_{D}$ in the introductory example as the sum instead of the average of all the instances. It should be noted that only 8 task instances of this potential can be realized by picking the better individual decision, while 5 task instances cannot be solved individually, but only—if at all—in collaboration. We reflect this by distinguishing two $CP$ components: inherent and collaborative complementarity potential.
Inherent complementarity potential represents improvements, i.e., loss reductions, that—from the perspective of the overall more accurate team member $T^{\ast}$—could be contributed by including any superior decisions on the instance level by the overall less accurate team member. Figure 2 illustrates this—assuming, without loss of generality, that AI is the overall better performing individual team member ($L_{AI}\leq L_{H}$). In respect of task instances where the inferior team member, i.e., the human, could help to reduce the loss, we capture the inherent complementarity potential.
Figure 2: Complementarity potential ($CP$) split into inherent ($CP^{inh}$) and collaborative ($CP^{coll}$) complementarity potential for a single instance with better human performance (left) and better AI performance (right) in respect of $T^{\ast}=AI$. $l_{D}$ denotes the instance-specific loss with $D\in\ \left\{H,AI\right\}$.
The remaining losses constitute the collaborative complementarity potential that could only be exploited if the team members’ collaboration yields new insights not available for the individual decisions before the collaboration.
Formally, inherent ${CP}^{inh}$ can be calculated by aggregating all the potential loss improvements for the overall better performing team member across all instances:
${CP}^{inh}=\begin{cases}\frac{1}{N}\sum\limits_{i=1}^{N}\max(0,l_{AI}^{(i)}-l_{H}^{(i)}),&L_{AI}\leq L_{H},\\ \frac{1}{N}\sum\limits_{i=1}^{N}\max(0,l_{H}^{(i)}-l_{AI}^{(i)}),&L_{AI}>L_{H}.\end{cases}$ (4)
The remaining collaborative ${CP}^{coll}$ can be calculated by aggregating the remaining minimum losses that the team members incur individually per task instance:
${CP}^{coll}=\frac{1}{N}\sum_{i=1}^{N}\min(l_{H}^{(i)},l_{AI}^{(i)})$ (5)
Collaborative potential signifies an improvement potential that goes beyond individual solutions in generating “new” knowledge. In respect of classification tasks, this could entail identifying the correct class decisions that neither match the (incorrect) human or AI decisions—based on the disagree of both. In respect of regression tasks, over- and underestimates of the respective team members may result in averaged estimates that are more accurate than the individual ones.
The inherent and collaborative components are additive and together form the complementarity potential (see Appendix A for additional details):
$CP={CP}^{inh}+{CP}^{coll}$ (6)
In our introductory example, the total complementarity potential of 13 instances can be differentiated into inherent complementarity potential (${CP}^{inh}$) of 8 instances (for which the human team member can contribut the correct solution), and a collaborative complementarity potential (${CP}^{coll}$) of 5, with none of the team members arriving at the correct decision individually.
Figure 3: Illustration of (theoretical) complementarity potential and (realized) complimentarity effect extending the introductory example.
3.4 Realization of Complementarity Potential
In real world collaboration scenarios between a human and AI, it is, of course unlikely that the entire complementarity potential will be exploited. We therefore introduce the complementarity effect ($CE$) as the part of this potential that is realized by the joint team decision. Measuring and dissecting this effect will allow observed human-AI team settings to be analyzed in greater detail, in order to infer conclusions about the collaboration’s effectiveness, and to purposefully develop and compare collaboration mechanisms. Analogous to the complementarity potential in Equation 3, the realized complementarity effect denotes the difference between the average loss of the overall individually better team member and that of the joint team decision ($L_{I}>0$):
$CE\ =\ min(L_{H},L_{AI})-L_{I}$ (7)
Figure 3 extends our introductory example from Figure 1 by additionally incorporating hypothetically realized human-AI decisions for all instances. We find that the human-AI team only makes 9 incorrect decisions compared to the 13 of the AI and the 15 of the human, when they act independently. This collaboration therefore reaches a complementarity effect of 4—realizing 31% of the full complementarity potential of 13.
Figure 4: Illustration of (theoretical) complementarity potential ($CP$) and (realized) complimentarity effect ($CE$) for different loss scenarios on an instance level—assuming, without loss of generality, that the AI performs better overall ($T^{\ast}=AI$). $l_{D}$ denotes the instance-specific loss with $D\in\ \left\{H,AI,I\right\}$.
The complementarity effect measures loss improvements—from the perspective of the overall more accurate team member $T^{\ast}$—that are realized by a particular combination of the individual decisions into a team decision. Analogous to inherent or collaborative potential, we can split the complementarity effect into the same two categories. Figure 4 illustrates this—again, we assume, without loss of generality, the AI to be the overall better performing individual team member ($L_{AI}\leq L_{H}$). If there is inherent potential (i.e., for task instances where the overall inferior human is more knowledgeable, as in scenario 1-3) this inherent complementarity potential may be realized partially ($l_{AI}>l_{I}>l_{H}$, scenario 1), fully ($l_{AI}>l_{H}>l_{I}$, scenario 2), or not at all ($l_{I}>l_{AI}$, scenario 3). In scenario 2, any improvement beyond the individual loss of the team member who is more accurate for a particular instance means that also the collaborative effect occurs. If there is no inherent complementarity effect (as in scenarios 3 and 4), any realization effect counts to the collaborative effect. We note, however, that the collaborative effect could also be negative, e.g., for $l_{I}>l_{AI}$ as in scenario 3. This would indicate that in these scenarios the collaboration rather deteriorates the outcomes. [^3] Note, that the negative collaborative effect should not be interpreted as unrealized inherent complementarity potential, as it occurs independently of the overall inferior team member’s individual loss.
In general, we can aggregate the complementarity effects across all instances of a task and summarize both cases (AI or human with overall better performance) and the scenarios above (depending on the instance performance of human, AI, and human-AI team):
${CE}^{inh}$ $=\frac{1}{N}\sum_{i=1}^{N}\begin{cases}l_{AI}^{(i)}-l_{I}^{(i)},&L_{AI}\leq L_{H}\text{ and }l_{AI}^{(i)}>l_{I}^{(i)}\geq l_{H}^{(i)},\\ l_{AI}^{(i)}-l_{H}^{(i)},&L_{AI}\leq L_{H}\text{ and }l_{AI}^{(i)}>l_{H}^{(i)}>l_{I}^{(i)},\\ l_{H}^{(i)}-l_{I}^{(i)},&L_{H}<L_{AI}\text{ and }l_{H}^{(i)}>l_{I}^{(i)}\geq l_{AI}^{(i)},\\ l_{H}^{(i)}-l_{AI}^{(i)},&L_{H}<L_{AI}\text{ and }l_{H}^{(i)}>l_{AI}^{(i)}>l_{I}^{(i)},\\ 0,&\text{otherwise}.\end{cases}$ (8)
${CE}^{coll}$ $=\frac{1}{N}\sum_{i=1}^{N}\begin{cases}l_{H}^{(i)}-l_{I}^{(i)},&l_{AI}^{(i)}\geq l_{H}^{(i)}>l_{I}^{(i)},\\ l_{AI}^{(i)}-l_{I}^{(i)},&l_{H}^{(i)}\geq l_{AI}^{(i)}>l_{I}^{(i)},\\ l_{AI}^{(i)}-l_{I}^{(i)},&L_{AI}\leq L_{H}\text{ and }l_{I}^{(i)}>l_{AI}^{(i)},\\ l_{H}^{(i)}-l_{I}^{(i)},&L_{H}<L_{AI}\text{ and }l_{I}^{(i)}>l_{H}^{(i)},\\ 0,&\text{otherwise}.\end{cases}$ (9)
Analogous to Equation 6, the inherent and collaborative components add up to the total complementarity effect (see Appendix A for additional details):
${CE=CE}^{inh}+{CE}^{coll}$ (10)
In our extended introductory example in Figure 3, we find that of the 8 task instances offering inherent complementarity potential, 3 could be realized as an inherent complementarity effect (${CE}^{inh}$) by taking the human’s individual decision suggestions into account. Of the 5 task instances for which neither the human nor the AI could individually make a correct decision, collaboration enabled the human-AI team to make correct decisions regarding 3 task instances. However, also 2 task instances that the AI alone could have performed correctly are subject to an erroneous decision due to the collaboration. Consequently, the collaborative complementarity effect (${CE}^{coll}$) amounts to 1 and the total complementarity effect ($CE$) to 4.
Figure 5 summarizes the complementarity potential, including its realization. It allows us to assess the collaboration between humans and AI more comprehensively. We will later illustrate the application of this concept in our experiments. Beforehand we, however, contemplate on the sources underlying the potential for complimentarity.
Figure 5: Overview of the complementarity potential and effect, including the inherent and collaborative components. Note that performance is represented by a loss that can be interpreted as an error measure. This means a lower value represents a better performance.
3.5 Sources of Complementarity Potential
A continuously growing body of research on human-AI decision-making assumes there are complementary capabilities between humans and AI (Bansal et al.,, 2021; Dellermann et al.,, 2019). The discourse is, however, held on a general level, e.g., by hypothesizing that humans excel at creativity, whereas AI better identifies patterns in large amounts of data (Dellermann et al.,, 2019). In this context, a conceptualization and empirical analysis of complementarity are still lacking. We therefore aim to provide a more nuanced understanding of complementarity potential by identifying and structuring its sources. To this end, we conceptually distinguish three phases in human-AI decision-making, as shown in Figure 6.
The first is the learning or training phase encompassing AI training and the human learning process. Note that AI models are usually developed within a short period of time through learning from aggregated training data, while human learning is a lifelong process. The second is the inference phase, in which the human and AI infer a decision pertaining to a particular instance. The third is the collaboration phase, in which the human and AI interact. Based on this simplified decision-making process, we discuss the sources of complementarity potential more granularly. We hypothesize that complementarity potential is created through asymmetries in either the information or the capabilities.
Training Phase.
First, training data as input of the training process differs between humans and AI. Humans might have experienced a considerable number of training instances, whereas AI is usually trained on a limited data set. Second, humans and AI’s inherent capabilities, which training stimulates, differ from one another. AI is, for instance, able to efficiently identify patterns in high-dimensional data or infer decisions from probabilistic reasoning (Dellermann et al.,, 2019; Jarrahi,, 2018). In contrast, humans are already able to learn abstract concepts from a small number of samples (Zheng et al.,, 2017). In this context, human decision-making is often purely heuristic or intuitive and does not consider all of the available information (Jarrahi,, 2018). Based on these asymmetries during the training process, humans and AI learn different decision boundaries from which their individual decisions are inferred (Geirhos et al.,, 2020, 2021; Steyvers et al.,, 2022). These differences could materialize in distinct capabilities; their presence results in inherent complementarity potential.
Figure 6: Conceptual overview of the human-AI decision-making process.
Inference Phase.
After the training process, the AI can infer a decision for a particular instance. Even when assuming that humans and AI have identical decision boundaries, different available information can constitute a source of inherent complementarity potential—in real-world settings, AI and humans often have access to different features (Bansal et al.,, 2019; Sanders and Ritzman,, 2001). A famous example is the “broken leg” scenario, which refers to side information known to humans (Meehl,, 1957). This information could not be incorporated as features into the model due to its rare occurrence. In many application domains in which AI is applied to support human decision-making, there is additional information beyond the data used to train the AI model (Ibrahim et al.,, 2021). In practice, due to technical or economic reasons, this discretionary data may often not be digitally available at all, or only in a small quantity insufficient for model training (Ibrahim et al.,, 2021). Nevertheless, in human-AI decision-making, human team members might leverage their expertise to use the additional information (Ibrahim et al.,, 2021). On the other hand, AI models could also have access to unique contextual information that is unavailable to human team members during the decision-making process.
To summarize, inherent complementarity potential can be present during training or inference and is usually either based on information or capabilities asymmetry.
Collaboration Phase.
The collaboration process adds a third phase to the overall decision-making process. In this phase, collaborative complementarity potential might arise as an additional part of complementarity potential beyond the inherent complementarity potential. For the collaborative aspect of complementarity potential, it is essential for the human-AI team to possess the capability to further improve the best individual decision achieved by either the AI or human in respect of a given task instance.
In the following, we illustrate the effects of information and capability asymmetry in behavioral experiments by applying the complimentarity concepts developed in this section.
4 Experimental Studies
To demonstrate the proposed conceptualization’s value and application, and to further investigate human-AI decision-making in the presence of the identified sources of complementarity potential—information and capability asymmetry—we conduct two behavioral experiments. Specifically, we focus on a team setting in which a human decision-maker has access to AI advice and is subsequently responsible for making the final team decision based on his/her judgment and that of the AI (Green and Chen,, 2019). In this collaboration setup, the team decision either matches the individual human or AI decision—i.e., the human relies fully on his/her judgment or that of the AI—or it can be a function of the individual human and AI decision—i.e., an “integrated” decision that potentially differs from both individual ones. In the first experiment, we investigate the effect of information asymmetry on decision-making in the human-AI team in the form of additional contextual information only available to the human. In the second experiment, we investigate the effect of capability asymmetry on joint decision-making with different levels of diversity between the capabilities of humans and AI.
4.1 Experiment 1: The Effect of Information Asymmetry
In the first experiment, we apply the conceptualization developed in Section 3 to study the effect of information asymmetry between humans and AI as a relevant source of complementarity potential. More precisely, we create an intervention in which humans are given contextual information withheld from the AI to investigate whether and how this affects the final team decision and the realized complementarity effect.
4.1.1 Task and AI Model
We draw on a real estate appraisal task provided on the data science website kaggle.com (Kaggle,, 2019). Since housing is a basic need, and because it is ubiquitous in everyone’s life, all people to some degree have the ability to assess a house’s value on the basis of relevant factors such as size or appearance. The data set encompasses 15,474 houses and contains information about the street, city, number of bedrooms, number of bathrooms, and size (in square feet). In the data set, the house prices denote their listing price. The average house price is \$703,120, ranging between a minimum of \$195,000 and a maximum of \$2,000,000. An image of each house is also provided. In respect of the house price prediction task, we implement a random forest regression model as the AI model (Breiman,, 2001). We draw on the individual trees in the random forest to generate a predictive distribution for each instance and provide the 5% and 95% quantiles as indicators of the AI’s prediction uncertainty. We use 80% of the data as the training set and 20% as the test set. We train the random forest on the following features: the street, city, number of bedrooms, number of bathrooms, and square feet of the house. The house’s image is withheld from the AI model. In respect of the behavioral experiment, we focus on detached family houses in the test set, with the existing image providing a view of its exterior. From these, we randomly draw a hold-out set of 15 houses to serve as samples for our behavioral experiment. The AI model achieves a performance measured in terms of the mean absolute error (MAE) of \$163,080 regarding the hold-out set, which is representative of its performance on the entire test set. In respect of the condition with unique human contextual information (UHCI), we give humans an additional image of the house, which is likely to constitute valuable information. Humans are able to leverage their general understanding to form an overall assessment based on the house’s features, the visible surroundings, and its appearance. We conduct an initial pilot study to verify this assumption (Appendix B contains additional details).
4.1.2 Study Design
We conduct an online experiment with a between-subjects design. We recruit participants from prolific.com. The study includes two conditions and randomly assigns each participant to one of these conditions. We do not allow any repeated participation. Each participant passes the following steps (see Appendix B for additional details on the study design):
Step 1: After accepting the task, participants are transferred to our experimental website. They are asked for their consent and to read the instructions. A control question initiates the study.
Step 2: Since prior work highlights the importance of task training for the participants, we include a mandatory tutorial (Grootswagers,, 2020). In order to familiarize the participants with the task and data, both treatments receive an identical in-depth introduction to the data set, including summary statistics like the mean and the maximum and minimum house prices as reference points. Participants in the treatment without unique human contextual information (i.e., without UHCI) are only given the houses’ tabular data, while the participants in the treatment with unique human contextual information (i.e., with UHCI) are also given images of the house.
Step 3: As part of the tutorial, we introduce the participants to the AI. We emphasize that the AI did not have access to the images during the training. We show the AI prediction in the context of the minimum and the maximum house prices. The participants also receive information about the AI’s uncertainty in the form of the 5% and 95% quantiles (see Figure 7). We explain the interpretation of the AI’s advice, including all the data points mentioned above. The participants are subsequently asked to answer a control question to verify their understanding.
Step 4: The participants conduct two training task instances. For each instance, we initially let the participants to first provide a prediction, before we reveal the AI’s recommendation. They are then asked to adjust the AI’s prediction in the best possible way, constituting the joint human-AI team prediction. After each training example, the participants receive feedback in the form of the true house price. After completing the two training task instances, they are informed about the start of the study.
Figure 7: An overview of the interfaces containing the information that the participants are given in the respective behavioral experiment’s treatments.
Step 5: Each participant completes 15 house price prediction task instances, presented randomly in the same procedure as described in Step 4. During the task, the participants are not informed about the actual house price. Subsequently, we ask them to complete a questionnaire regarding qualitative feedback (Step 6) and demographic information (Step 7).
The overall task lasts approximately 30 minutes. Before recruiting participants, we compute the required sample size in a power analysis using G*Power (Faul et al.,, 2007). Based on the pilot data, we expect a large effect ($d=$ 0.8). We refer to an alpha value of 0.05, while taking multiple testing into account in order to achieve a power of 0.8. This results in a total sample size of 86. Anticipating that some participants will fail the attention checks, we recruit a total of 120 participants (60 per condition). They receive a base payment of £5 and are additionally incentivized following the approach of Kvaløy et al., (2015), who show the benefits of combining non-monetary motivators, such as recognition, attention, and verbal feedback, with performance-based pay. We achieve this by adding motivational statements and by giving the top 10% participants an additional pound. Note that the two training task instances are not included in the final evaluation. To ensure the quality of the collected data, we remove those participants whose entered prices exceed the communicated maximum house price of \$2,000,000 in the data set. We also identify outliers for removal by using the median absolute deviation (Leys et al.,, 2013; Rousseeuw and Croux,, 1993). After applying these criteria, we continue with the data of 101 participants across both conditions—53 in the treatment without UHCI and 48 in that with UHCI (see Appendix B for additional details about the participants).
4.1.3 Evaluation Measures
For each participant, we measure the loss of the human ($l_{H}$), the AI ($l_{AI}$), and the team decision ($l_{I}$) as the absolute error, and calculate the average over all task instances to receive the human ($L_{H}$), the AI ($L_{AI}$), and the team performance ($L_{I}$) corresponding to the mean absolute error (MAE). Furthermore, we calculate the complementarity potential’s and complementarity effect’s respective components as defined in Section 3. Finally, for each measure, we calculate the average over all the participants.
4.1.4 Results
In this section, we analyze the impact of unique human contextual information on the team performance, the complementarity potential, and the complementarity effect. We evaluate the results’ significance by using the Student’s T-test and the Mann-Whitney U-test, depending on whether the prerequisites have been fulfilled. We apply the Bonferroni correction and adjust the p-values accordingly. First, we focus on the impact of contextual information on performance, followed by an in-depth analysis of its effect on the complementarity potential and on its constituting components.
Figure 8 displays the isolated human and joint human-AI performance for both conditions. It also includes the performance of the AI alone. We first evaluate the impact of unique human contextual information without any AI assistance. Participants in the treatment without UHCI achieve an MAE of \$251,282, while those in the treatment with UHCI yield an MAE of \$200,510—an improvement of \$50,772 (20.21%), which is significant ($d=$ 0.92, $p<$ 0.001, two-sample, two-tailed T-test). This result confirms the general usefulness of the provided house images from the human perspective.
Next, we evaluate the impact of unique human contextual information when the human is teamed with the AI. The team performance in the treatment without UHCI results in an MAE of \$160,095 versus an MAE of \$148,009 in the treatment with UHCI—an improvement of \$12,086 (7.55%), which is significant ($d=$ 0.59, $p<$ 0.05, two-sample, two-tailed T-test). In both treatments, the human-AI team outperforms the AI (MAE: \$163,080). Whereas the difference between the performance of the human-AI team and the performance of the AI alone is significant in the treatment with UHCI ($d=$ 0.68, $p<$ 0.001, one-sample, two-tailed T-test), the difference in the treatment without UHCI does not constitute a significant improvement ($d=$ 0.16, $p=$ 1.0, one-sample, two-tailed T-test).
Figure 8: Performance results as the MAE across the conditions (UHCI $=$ unique human contextual information), including 95% confidence intervals. The red horizontal line denotes the AI performance.
Complementarity Potential.
First, we analyze the inherent complementarity potential (${CP}^{inh}$). We observe a significant increase due to the unique human contextual information. In the condition without UHCI, the ${CP}^{inh}$ is \$42,995, and increases to \$61,970 in the condition with UHCI ($d=$ 1.05, $p<$ 0.001, two-tailed Mann-Whitney U test). This finding can be interpreted that the images indeed contain useful contextual information for humans, which the AI cannot access.
Next, we calculate the collaborative complementarity potential (${CP}^{coll}$). Whereas in the condition without UHCI, the ${CP}^{coll}$ results in \$120,085, in the condition with UHCI it amounts to \$101,110. The difference is statistically significant ($d=$ 1.05, $p<$ 0.001, two-tailed Mann-Whitney U test). Finally, since the participants in both conditions work with the same AI, which has an overall better individual performance, the $CP$ (i.e., the sum of the inherent and collaborative component) is \$163,080 in both conditions. As ${CP}^{inh}$ increases due to UHCI, ${CP}^{coll}$ decreases because the $CP$ remains constant.
Complementarity Effect.
Next, we focus on the realized complementarity potential, i.e., the complementarity effect ($CE$). We find a significant difference between the inherent complementarity effect (${CE}^{inh}$) in both conditions (without UHCI: \$14,468; with UHCI: \$27,860; $d=$ 0.87, $p<$ 0.001, two-tailed Mann-Whitney U test), which highlights contextual information’s potential. This absolute increase might be due to an increase in the complementarity potential and/or an improvement in the integration of both team members’ predictions through the human. In this context, in order to investigate this further, we also calculate the inherent complementarity effect’s ($\frac{{CE}^{inh}}{{CP}^{inh}}$) relative amount. This analysis reveals that unique human contextual information not only enhances the theoretically available inherent complementarity potential, but that the participants could also use significantly more of it (without UHCI: 34%; with UHCI: 45%; $d=$ 0.82, $p<$ 0.001, two-tailed Mann-Whitney U test). This is an interesting result, as having more information available than the AI might also have detrimental psychological effects. Jussupow et al., (2020), for example, find that perceived AI capabilities and expertise influence aversion towards AI. This could lead to humans taking less account of the AI’s suggestions when making their final decision (Longoni et al.,, 2019; Mahmud et al.,, 2022; Sieck and Arkes,, 2005), which could result in $CTP$ not being achieved. However, our results show that unique human contextual information does not only increase the potential, but also the integration’s overall effectiveness.
Figure 9: Summary of the real estate appraisal experiment’s results.
Next, we analyze unique human contextual information’s impact on the collaborative complementarity effect (${CE}^{coll}$). We do not find a significant difference between the two treatments (${CE}^{coll}$: without UHCI: \$-11,483; with UHCI: \$-12,789; $d=$ 0.08, $p=$ 1.0, two-tailed Mann-Whitney U test). In our experiment, it seems intuitive for the ${CE}^{coll}$ to yield a negative value. Given that the AI in our setup outperforms its human team member, a positive ${CE}^{coll}$ could only occur on task instances where the team loss is even lower than that of the AI and the human (see Figure 4). Conversely, each instance where the human underperforms in comparison to the AI and fails to adopt the AI’s decision contributes negatively to the ${CE}^{coll}$. Since humans tend to choose decisions between two boundaries (in our case their own and the AI’s decision), our setup naturally fosters a negatively realized collaborative complementarity potential. See Appendix B for further results, including a performance analysis of each house.
Overall, the most important finding is that humans can realize a disproportionally large amount of the inherent complementarity potential through unique contextual information, which finally results in $CTP$. This constitutes a new empirical insight that could only be measured due to the granular formalization. Summing ${CE}^{inh}$ and ${CE}^{coll}$ results in the total complementarity effect ($CE$), which equals the performance difference between the best individual team member and the joint human-AI team performance (without UHCI: \$2,985; with UHCI: \$15,071). Figure 9 summarizes the results of our experiment.
4.2 Experiment 2: The Effect of Capability Asymmetry
In the second behavioral experiment, we apply the conceptualization to study the effect of capability asymmetry between humans and AI as another relevant source of complementarity potential. In detail, starting with a “baseline” AI, we create an intervention in which we increase the asymmetry between the AI’s and the humans’ capabilities while keeping its overall performance constant. This means the second AI tends to make correct decision suggestions for task instances that tend to be more difficult for humans (“complementary” AI).
4.2.1 Task and AI Model
To investigate the effect of capability asymmetry on the human-AI team performance in decision-making, we choose the image recognition context. Research has demonstrated that humans and AI tend to make different errors on image classification tasks (Fügener et al., 2021a, ; Steyvers et al.,, 2022). Specifically, an AI based on deep convolutional neural networks tends to infer classification decisions differently than humans do (Geirhos et al.,, 2020, 2021). We could therefore expect the AI model to classify certain images more accurately than humans and vice versa, thereby creating inherent complementarity potential. However, it remains unclear whether this naturally existing potential could be sufficiently realized when humans incorporate the AI decision into a final team decision, and whether its increase in the intervention affects the realization.
In order to undertake the experiment, we draw on the image data set that Steyvers et al., (2022) provided. The data set comprises 1,200 images distributed evenly across 16 classes (e.g., airplane, dog, or car). It was curated on the basis of the ImageNet Large Scale Visual Recognition Challenge (ILSVRC) 2012 database (Russakovsky et al.,, 2015). To increase the task difficulty for humans and the AI, the authors applied phase noise distortion at each spatial frequency, which was uniformly distributed in the interval $[-\omega,\omega]$ with $\omega=$ 110. Despite the heightened difficulty level, both humans and AI can attain comparable performance on the task. In addition to ground truth labels, the data set also contains multiple human predictions for each image provided by crowd workers, allowing us to infer a proxy for human classification difficulty. Images with a high disagreement in respect of multiple human predictions indicate a higher level of difficulty for humans.
We implement the AI model as a convolutional neural network, more precisely, as a DenseNet161 (Huang et al.,, 2017), pre-trained on ImageNet. We partition the data set into a training (60%), validation (20%), and test set (20%). In the baseline condition, we fine-tune the AI model on the distorted images over 100 epochs, applying early stopping on the validation loss. We use SGD as an optimizer with a learning rate of $1\cdot{10}^{-4}$, a weight decay of $5\cdot{10}^{-4}$, a cosine annealing learning rate scheduler, and a batch size of 32. The AI model achieves a classification error of 26.66% on the test set.
In the intervention, we create an alternative AI model that makes erroneous decisions for different instances. We fine-tune the DenseNet161 model exactly as in the baseline condition, but, for each image in the training set, also incorporate a human prediction as an additional label in the training process in order to incentivize the AI model to learn to correctly classify the images that tend to be more difficult for humans (Hemmer et al.,, 2022; Madras et al.,, 2018; Wilder et al.,, 2020). See Appendix C for additional implementation details of this approach. Even though the AI model has a slightly higher classification error of 33.75%, this approach results in more asymmetric, i.e. non-overlapping, capabilities between the humans and the AI model. We select 15 images from the test set for the experiment, such that both AI models exhibit the same performance on the test set (26.66%), while considering non-overlapping errors between both AI models.
4.2.2 Study Design
We conduct an online experiment, employing a between-subjects design with two conditions. Participants are recruited from prolific.com and randomly assigned to one of the conditions; repeated participation is not allowed. We employ a similar experimental set up as in the first experiment (see Appendix C for additional details on the study design):
Step 1: The participants are transferred to the experimental website after accepting the task, after which they have to submit a consent form and answer an initial control question. Thereafter they have to pass an attention check.
Step 2: They are given a task tutorial which presents an exemplary image along with the 16 classes arranged in a four-by-four matrix, including the 16 class icons with their name displayed underneath.
Step 3: The participants are given an introduction to the AI and its decision regarding an exemplary image followed by a control question to verify their understanding. Figure 10 displays the interface that the participants see throughout the experiment when collaborating with the AI. In this experiment, the participants are only shown the AI’s decision, since confidence information generated by two separately trained AI models would introduce a confounder.
Figure 10: An overview of the interfaces that the participants are shown in both treatments of the behavioral experiment.
Step 4: Next, participants start a practice round comprising three images that need to be classified without AI support in order to familiarize themselves with the classification task.
Step 5: After being informed about the start of the main task, each participant classifies 15 images in a randomized order. The participants provide their own classification for each image, then receive the AI recommendation and are asked to verify and adjust, if required, the AI classification in the best way possible. During the task, they are not informed about the true class of any image. After classifying all the images, participants are asked to answer a questionnaire regarding qualitative feedback (step 6) and demographic information (step 7).
The overall task lasts approximately 20 minutes. Before recruiting participants, we compute the required sample size in a power analysis using G*Power (Faul et al.,, 2007). We test for a medium to large effect ($d=$ 0.65) and refer to an alpha value of 0.05, while taking multiple testing into account in order to achieve a power of 0.8. This results in a total sample size of 128. In order to buffer for participants potentially failing attention checks, we recruit a total of 170 participants. They receive a base payment of £8 and are additionally incentivized following the approach pursued in the first behavioral experiment (Kvaløy et al.,, 2015). We exclude participants who did not pass the integrated attention checks, resulting in 144 participants—76 in the base condition (baseline AI) and 68 in the intervention (complementary AI). We provide further details about the participants in Appendix C.
4.2.3 Evaluation Measures
For each participant, the loss of the human ($l_{H}$), the AI ($l_{AI}$), and the team decision ($l_{I}$) is measured as the classification error and averaged over all the task instances, providing the human ($L_{H}$), the AI ($L_{AI}$), and the team performance ($L_{I}$). We also calculate the complementarity potential and effect, including their components (see Section 3). Finally, for each measure, we calculate the average over all the participants.
4.2.4 Results
We analyze the effect of capability asymmetry on team performance, complementarity potential, and complementarity effect while assessing the statistical significance by using the same procedure as in the first experiment.
Figure 11 shows the classification error for humans performing the task alone and together with the AI in both conditions. In addition, it also includes the classification error of both AI models, which are identical in this task. Humans conducting the task alone exhibit a classification error of approximately 0.30, which is nearly identical across the conditions (Baseline AI: 0.2999; Complementary AI: 0.2951; $d=$ 0.05, $p=$ 1.0, two-sample, two-tailed T-test). When humans are teamed with the AI, the joint performance increases in both conditions. Whereas the human-AI team yields a classification error of 0.2473 in the condition with the baseline AI, this error decreases even further to 0.1461 in the team with the complementary AI. This corresponds to an improvement of 41%, which is significant ($d=$ 1.29, $p<$ 0.001, two-sample, two-tailed T-test). Both classification errors are significantly lower than that of the AI conducting the task alone in both conditions (Baseline AI: 0.2666, $d=$ 0.33, $p<$ 0.05, one-sample, two-tailed T-test; Complementary AI: 0.2666, $d=$ 1.25, $p<$ 0.001, one-sample, two-tailed T-test).
Figure 11: Performance results as the classification error across the conditions, including 95% confidence intervals. The red horizontal line denotes the AI performance.
Complementarity Potential.
We observe a significant increase in the inherent complementarity potential (${CP}^{inh}$) in the condition with the complementary AI (Baseline AI: 0.0640, Complementary AI: 0.2480; $d=$ 3.81, $p<$ 0.001, two-tailed Mann-Whitney U test). This reflects the AI model’s more asymmetric capabilities compared to that of the humans in this condition. Compared to the baseline AI condition, the AI model makes more erroneous decisions for instances that humans can process correctly, whereas humans err for instances that AI can classify correctly. Conversely, we observe a significant decrease in the collaborative complementarity potential (${CP}^{coll}$) in the condition with the complementary AI (Baseline AI: 0.2026, Complementary AI: 0.0186, $d=$ 3.81, $p<$ 0.001, two-tailed Mann-Whitney U test) as the $CP$ remains constant due to the AI being individually more accurate than the humans. Whereas the inherent complementarity potential constitutes 24% of the overall complementarity potential in the baseline condition, this share rises to 93% in the complementary AI condition. This means that, for the majority of the task instances, one team member is theoretically capable of making a correct decision, provided that the human responsible for the team decision relies on his/her own assessment or that of the AI in respect of the correct instances.
Complementarity Effect.
In the baseline condition, 58% of the inherent complementarity potential ($\frac{{CE}^{inh}}{{CP}^{inh}}$) could be realized by the humans integrating their own decision and that of the AI into a final team decision, resulting in a ${CE}^{inh}$ of 0.0368. In the condition with the complementary AI, it is possible to realize 89% of the inherent complementarity potential, resulting in a ${CE}^{inh}$ of 0.2196. This not only means a significant performance improvement ($d=$ 3.43, $p<$ 0.001, two-tailed Mann-Whitney U test), but also that a significantly larger fraction of $\frac{{CE}^{inh}}{{CP}^{inh}}$ could be realized ($d=$ 1.02, $p<$ 0.001, two-tailed Mann-Whitney U test). It indicates that humans tended to rely on the AI decisions when they were correct, but on their decision when it was incorrect. Moreover, in both conditions the collaborative complementarity effect (${CE}^{coll}$) is negative. Whereas the value is only slightly negative in the baseline condition (Baseline AI: -0.0175), it decreases to -0.0990 in the condition with the complementary AI. The difference between the two conditions is significant ($d=$ 1.32, $p<$ 0.001, two-tailed Mann-Whitney U test).
Figure 12: Summary of the results of the image classification experiment.
The observed changes in both of the components reveal interesting insights. The increase in the inherent complementarity effect reveals that humans continue to regard the AI as a team member on whom they can rely, especially with regard to the instances that they find difficult, even though they have observed the AI makes several incorrect decisions in respect of the instances that they find easier. Conversely, the decrease in the collaborative complementarity effect reveals that there are also instances where humans, incorrectly, do not rely on the AI. This is an interesting finding, since it could have been conceivable for humans to avoid AI suggestions to a greater extent after witnessing incorrect AI decisions (Dietvorst et al.,, 2015). Consequently, in the team setting in which the human integrates his or her own decision and that of the AI into a final team decision, there might be a trade-off between turning the AI’s more asymmetric capabilities into performance synergies and not relying on its suggestions due to witnessing erroneous decisions in “easier” instances. See Appendix C for additional results, including a performance and a switch fraction analysis of each image. In summary, ${CE}^{inh}$ and ${CE}^{coll}$ result in the total complementarity effect ($CE$)—equivalent to the performance difference between the best individual team member and the team performance (Baseline AI: 0.0193, Complementary AI: 0.1206). Figure 12 summarizes the analysis.
5 Discussion
Our research emphasizes the importance of considering $CTP$ in human-AI decision-making. In the past, when early AI-based systems still had limited capabilities, $CTP$ was rarely a focus, as these systems could only provide humans with (partial) decision support within a small range of low-stakes decision-making tasks (e.g., calculating decision-making inputs) (Turban et al.,, 2011). Today, however, AI can perform a growing number of tasks independently with human-level performance, and is increasingly utilized in high-stakes decision-making domains such as medicine (McKinney et al.,, 2020), law (Hillman,, 2019), and finance (Day et al.,, 2018). It can also provide general-purpose support, driven by recent advances in large language models that enable applications such as ChatGPT (Bubeck et al.,, 2023). While this opens up the potential for task automation, it also enables a new form of human-AI decision-making with AI models and humans being equitable team members.
5.1 Contributions
Current research lacks a concise concept of human-AI complementarity, as the majority of studies focusing on human-AI decision-making do not to achieve $CTP$ (Bansal et al.,, 2021; Hemmer et al.,, 2021). Moreover, merely observing whether $CTP$ has been achieved or not, does not allow in-depth conclusions to be drawn about the synergetic potential of humans and AI through collaboration. In this work, we contribute to a comprehensive understanding of human-AI teams’ inner workings regarding decision-making, and provide guidance on how to achieve $CTP$ more consistently.
More specifically, we develop a conceptualization of human-AI complementarity that introduces the notion of complementarity potential in a formalization and delineates relevant sources. This conceptualization allows for analyzing human-AI teams’ potential for decision-making synergies and for making this potential measurable. By differentiating between the inherent and collaborative components of complementarity potential and its realization, it is also possible to gain important insights into the collaboration’s functioning in terms of the optimal joint team decision. These insights, along with the sources of complementarity potential, could inform human-AI teams’ future designs and collaboration mechanisms, depending on the use case and application domain. Furthermore, we not only demonstrate the conceptualization’s utility in two behavioral experiments with humans in the role of integrating the final team decision, but we also empirically show that both information and capability asymmetry can lead to $CTP$. In this context, our conceptualization allows us to reveal interesting empirical insights. The first experiment highlights that providing humans with unique contextual information not only affects the inherent complementarity potential, but can also increase the realized amount, i.e., the inherent complementarity effect, disproportionally. This means that the effectiveness of the integration that humans conducted improved, which is an interesting finding. Intuitively, the perception of having more information than the other team member could alternatively result in reduced utilization of the AI suggestions in the team prediction—resulting in a sub-optimal team performance (Jussupow et al.,, 2020; Mahmud et al.,, 2022; Sieck and Arkes,, 2005). The second experiment reveals that if the human-AI team’s capability asymmetry is greater, this can contribute to the realization of a larger share of the existing complementarity potential. This is also an insightful observation, since humans observing erroneous AI decisions for instances that they could solve relatively easily themselves, might lead them to refute AI support even in situations when it is actually helpful (Dietvorst et al.,, 2015). This finding also contributes to our understanding of team diversity’s compositional impact on performance (Horwitz,, 2005). Whereas expertise diversity in human teams could be a performance driver, e.g., because it fosters a broader range of cognitive skills (Cohen and Bailey,, 1997), it could also turn into a performance inhibitor, e.g., due to the potential difficulties of achieving a mutual understanding (Dougherty,, 1992). As in human teams, there might also be a trade-off, since humans could start avoiding AI suggestions if they observe them erring too often (Dietvorst et al.,, 2015). Our conceptualization enables an analysis of this tradeoff—finding that, in the specific experiment, the capability asymmetry between humans and AI affects team performance positively.
5.2 Theoretical Implications
From a theoretical perspective, the proposed conceptualization provides a foundation for future research in human-AI decision-making. We offer the research community concrete measures that allow the investigation of human-AI decision-making’s inner workings at a deeper level than aggregate performance comparisons in behavioral experiments allow. In this context, these measures could also help researchers formulate and test hypotheses about the effect of different socio-technical factors on the realization of decision-making synergies in human-AI teams (Jain et al.,, 2021). Our research’s most important implication is the need to design for $CTP$, which is influenced by the complementarity potential and the collaboration mechanism to derive a team decision. Both could and should be purposefully designed. The inherent complementarity potential can be influenced by increasing unique knowledge. From an AI perspective, this could be achieved by deliberately creating “complementary” AI models designed for team work that perform well in areas of the feature space where humans experience difficulties (Hemmer et al.,, 2022; Mozannar and Sontag,, 2020; Wilder et al.,, 2020). From a human perspective, humans could be trained to focus on their unique capabilities and to build awareness to use unique contextual information. Lastly, realizing complementarity potential also implies that the collaboration mechanism should be consciously designed to optimize the realization of inherent and collaborative complementarity potential.
5.3 Managerial Implications
Our work also has important implications for managerial decision-makers. In application areas suitable for human-AI decision-making tasks, responsible managers should focus on deploying AI systems that enable the realization of $CTP$. If they don’t, their competitors could gain advantages. They could start by collecting data to train AI models that are compatible with team work and to invest in trainings to upskill their employees in order to enhance their human-AI teams’ inherent complementarity potential. The conceptualization supports the identification of use cases suitable for human-AI decision-making and their evaluation of $CTP$’s realization. Current AI endeavors often result in the blind adoption of human-AI collaboration due to decision-makers’ concerns about full automation (Jussupow et al.,, 2021). However, it is worth designing human-AI collaboration purposefully, as this could contribute to a reduction in erroneous decisions with potentially high costs. Rather than fearing automation, decision-makers should explore the benefits of working with AI. This is an important perspective, also in the broader societal debate on the role of AI in automation and in the future of work. Our research provides managers with valuable guidance by helping them determine when and how to collaborate with AI to optimize decision-making and achieve $CTP$.
5.4 Limitations
Our current research has several limitations that future work needs to address. First, the prevalent use of laboratory-based experiments and vignette methodologies in extant literature on human-AI collaboration, including our own, could limit our work’s generalizability and its practical implications. This underscores the need for future studies to use a nuanced approach to prevent contributing to the fragmentation of research in this domain. Furthermore, we focus on developing measurement tools, delineating sources of complementarity potential, and on validating the proposed concepts in behavioral experiments. While our experiments’ controlled settings allow us to derive the insights presented in this work, they do not yet address the wide range of complex human factors, such as motivation (Schunk,, 1995), engagement (Chandra et al.,, 2022), or behavior patterns within teams (Schecter et al.,, 2022), which also contribute to human-AI collaboration’s effectiveness (Chandra et al.,, 2022). In this context, many factors in socio-technical systems are highly interconnected and influence the underlying systems’ acceptance and use (Jain et al.,, 2021). While a broader exploration of these factors is beyond the scope of our work, advancing the field of human-AI collaboration further also requires their incorporation. We hope that our research will serve as a foundation of studies aiming at expanding knowledge of the mechanisms governing human-AI collaboration.
Another limitation is the way in which we measure the counterfactual human decision if the user has not received advice on using AI. In this work, we used a sequential decision-making setup to first measure the human decision and, thereafter, the team decision. However, the sequential nature of the decision-making process could also influence human behavior. Consequently, the timing when the AI’s recommendation is revealed is another critical aspect (Jussupow et al.,, 2021). Giving the participants the AI’s recommendation upfront, could lead to cognitive capacity being less invested in the task, because the AI has already provided a possible answer (Green and Chen,, 2019).
5.5 Future Work
There are several potential areas of future research on human-AI complementarity. The conceptualization could obviously be applied to other domains in the future, but there are also methodological avenues to pursue.
Future work should expand our knowledge about relevant sources of complementarity potential. In this work, we have shown that information and capability asymmetry could be promising sources of complementarity potential. Regarding information asymmetry, we experimentally evaluated unique human contextual information. Investigating unique AI contextual information in future work could also produce interesting insights. It would, moreover, be worthwhile developing criteria regarding the utility of contextual information and the degree of capability asymmetry.
Furthermore, team settings could comprise more than two team members, considering multiple AI models or multiple humans. Team design principles could be derived on how to ensure a sufficient amount of complementarity potential and on how to select and combine human and artificial team members.
Different collaboration mechanisms could also be developed and evaluated. While we have focused on humans handling the integration in both experimental studies, the AI might conceivably also undertake the integration. Alternatively, decisions might not be integrated, but task instances might be delegated to a team member deciding on behalf of the team, with the delegation initiated by either the human or the AI (Fügener et al.,, 2022). Future work could investigate these collaboration forms in terms of the realization of complementarity potential and their suitability for specific use cases and application domains.
6 Conclusion
So far, human-AI decision-making has been primarily concerned with AI systems helping human users. However, since the number of decision tasks that can be automated (i.e., can be solved by the AI alone) is increasing steadily, the focus has shifted to the purposeful design of the collaboration between humans and AI as team members—thereby shaping the future of work with AI. The ultimate objective of these teams must be to achieve complementary team performance ($CTP$), with the team outperforming each individual team member. The IS community is predestined to drive the development of appropriate theories and to lay the foundation for practical applications. We hope that the conceptual foundation developed in this paper will provide fruitful ground for future research, and that the empirical studies illustrate the validity and potential of the human-AI complementarity paradigm.